# Styl3R: Instant 3D Stylized Reconstruction for Arbitrary Scenes and Styles

### v1

@@ 3.3 Training Curriculum
=$\mathcal{L}=\begin{cases}\mathcal{L}_{\text{photo}}(\mathcal{G}^{c}),&\text{NVS pre-training}\\ \lambda\mathcal{L}_{\text{style}}(\mathcal{G}^{s})+\mathcal{L}_{\text{content}}(\mathcal{G}^{s})+\mathcal{L}_{\text{photo}}(\mathcal{G}^{c}),&\text{stylization fine-tuning}\end{cases}$ (2)
@@ 4 Experiments
Method Consistency Stylization Time
@@
[…]
2D AdaIN [11] 0.163 0.063 0.323 0.111 0.004 s
@@
Because of the novel and under-explored nature of 3D stylization, there are few metrics for assessing the quality of the stylization. Therefore, we evaluate the multi-view consistency as in prior 3D stylization works [7, 23, 22]. We estimate optical flow between sequential images using RAFT [40], then warp the earlier frame with softmax splatting [28]. Consistency is measured by LPIPS [52] and RMSE between the warped and target images over valid pixels. Short- and long-range consistency are computed between adjacent views and those seven frames apart, respectively. To evaluate novel view synthesis quality, we report standard image similarity metrics: PSNR, SSIM, and LPIPS [52].
@@
We use PyTorch. The content and style encoder adopts a standard ViT-Large architecture with a patch size 16, while the structure and stylization decoder is based on a ViT-Base model. We initialize the encoder, decoder, and the Gaussian center prediction head with pretrained weights from MASt3R [18], whereas the remaining layers are initialized randomly. The model is trained on images with a resolution of $256\times 256$ pixels. Besides, we use 0 degree spherical harmonics for Gaussians following [7]. Training takes ~1.5 days on 8 NVIDIA A100 GPUs.
@@ 4.1 Experimental Results
As shown in Fig. 3 and Table 2, our method significantly outperforms all baselines. Visually, our stylizations achieve a more balanced trade-off between content preservation and faithful style transfer. Among test-time optimization-based 3D baselines, StyleRF [22] and StyleGaussian [23] often fail to reproduce the reference style’s color tone accurately, resulting in overly whitened or darkened outputs. ARF [51], while better at capturing style colors, tends to overfit and apply excessive stylization that obscures scene details. For example, in the third row of Fig. 3, furniture in the living room becomes nearly indistinguishable due to heavy sketch-line artifacts. As a 2D baseline, StyTr2 [4] generates visually pleasing results on individual ground-truth novel views, but it lacks multi-view consistency, as evidenced in Table 2 and Fig. 5. In contrast, our method consistently produces superior stylizations while maintaining the best short- and long-range consistency metrics, benefiting from our attention mechanism that operates jointly over multi-view content and style tokens. Although StyleRF achieves a slightly lower RMSE in short-range evaluations, this is largely due to its over-smoothed outputs, as clearly illustrated in Fig. 3. Video results in supplementary.
@@ 4.2 Ablation Studies
As discussed in Sec. 3.3, using both relu3_1 and relu4_1 for the content loss better preserves structural details without sacrificing artistic expression. As shown in Fig. 7, relying solely on relu3_1 tends to cause the style to overwhelm the underlying scene structure.
@@ 5 Conclusion
  • more visual results of our model and comparisons with baselines in Sec. B.

### v2

@@ 3.3 Training Curriculum
$\mathcal{L}=\begin{cases}\mathcal{L}_{\text{photo}}(\mathcal{G}^{c}),&\text{NVS pre-training}\\ \lambda\mathcal{L}_{\text{style}}(\mathcal{G}^{s})+\mathcal{L}_{\text{content}}(\mathcal{G}^{s})+\mathcal{L}_{\text{photo}}(\mathcal{G}^{c}),&\text{stylization fine-tuning}\end{cases}$ (2)
@@ 4 Experiments
Method Consistency
Stylization
Time
@@
$downarrow LPIPS\downarrow RMSE\downarrow$
2D
AdaIN [12] 0.163 0.063 0.323 0.111 0.004 s
@@
Because of the novel and under-explored nature of 3D stylization, there are few metrics for assessing the quality of the stylization. Therefore, we evaluate the multi-view consistency as in prior 3D stylization works [8, 24, 23]. We estimate optical flow between sequential images using RAFT [41], then warp the earlier frame with softmax splatting [29]. Consistency is measured by LPIPS [55] and RMSE between the warped and target images over valid pixels. Short- and long-range consistency are computed between adjacent views and those seven frames apart, respectively. We further employed ArtFID [46], a metric well aligned with human perceptual judgment by jointly assessing content preservation and style fidelity, together with the RGB-uv histogram from HistoGAN [1] to comprehensively evaluate the quality of color transfer. To evaluate novel view synthesis quality, we report standard image similarity metrics: PSNR, SSIM, and LPIPS [55].
Styl3R ARF StyleGS StyleRF StyTr2
RE10K (%) 53.29 14.13 6.67 1.81 24.10
TnT (%) 38.78 22.11 11.56 3.06 24.49
Table 3: User Study. Voting results for stylization.
Metric Styl3R ARF StyleGS StyleRF StyTr2
[…]
Table 4: ArtFID [46] and Histogram [1] for comparisions.
ArtFID $\downarrow$ Histogram $\downarrow$ Votes (%)
h3 layer 42.01 0.277 35.20
h3+h4 layer 35.83 0.239 64.80
Table 5: Ablations. The ArtFID, Histogram, and voting results for h3 and h3+h4 layer.
User Study.
We conducted an online user study to evaluate perceptual preferences between our proposed approach and alternative methods. Participants were asked to select the result they perceived as exhibiting the most harmonious integration of style and scene. In total, approximately 2,000 votes were collected, including 1,274 for RE10K, 294 for Tanks and Temples, and 392 for the ablation study.
@@
We use PyTorch. The content and style encoder adopts a standard ViT-Large architecture with a patch size 16, while the structure and stylization decoder is based on a ViT-Base model. We initialize the encoder, decoder, and the Gaussian center prediction head with pretrained weights from MASt3R [19], whereas the remaining layers are initialized randomly. The model is trained on images with a resolution of $256\times 256$. Besides, we use 0 degree spherical harmonics for Gaussians following [8]. Training takes ~1.5 days on 8 NVIDIA A100 GPUs.
@@ 4.1 Experimental Results
As shown in Fig. 3,  Table 2, Table 5 and Table 5, our method outperforms all baselines qualitatively and quantitatively. Visually, our stylizations achieve a more balanced trade-off between content preservation and faithful style transfer. Among test-time optimization-based 3D baselines, StyleRF [23] and StyleGaussian [24] often fail to reproduce the reference style’s color tone accurately, resulting in overly whitened or darkened outputs. ARF [54], while better at capturing style colors, tends to overfit and apply excessive stylization that obscures scene details. For example, in the third row of Fig. 3, furniture in the living room becomes nearly indistinguishable due to heavy sketch-line artifacts. As a 2D baseline, StyTr2 [5] generates visually pleasing results on individual ground-truth novel views, but it lacks multi-view consistency, as evidenced in Table 2 and Fig. 5. In contrast, our method consistently produces superior stylizations while maintaining the best short- and long-range consistency metrics, benefiting from our attention mechanism that operates jointly over multi-view content and style tokens. Although StyleRF achieves a slightly lower RMSE in short-range evaluations, this is largely due to its over-smoothed outputs, as clearly illustrated in Fig. 3.
@@ 4.2 Ablation Studies
As in Sec. 3.3, using both relu3_1 and relu4_1 for the content loss better preserves structural details without sacrificing artistic expression. As shown in Fig. 7 and  Table 5, relying solely on relu3_1 tends to cause the style to overwhelm the underlying scene structure.
@@ 4.3 Application
5 Discussion
To clarify the motivation of our method, we provide a detailed comparison between two-stage approaches, which first reconstruct 3DGS and then perform stylization, and our single-pass method that achieves 3D stylization in an end-to-end manner. [^1] Throughout this section, “(+)” and “(–)” denote advantages and disadvantages, respectively.
Two-stage methods:
  • (+) High-fidelity geometry: The 3DGS model is optimized directly to fit the target scene from dense image inputs, resulting in accurate geometry and detailed structures.
  • (+) Scalable to large scenes: With sufficient optimization time and resources, this approach can be extended to larger-scale environments.
  • (–) No instant stylization support: There is currently no method that directly takes a trained 3DGS and outputs a stylized version without additional test-time optimization. Even the closest work (e.g., StyleGaussian) requires post-processing like VGG feature embedding.
  • (–) Inefficient: The optimization-heavy pipeline is unsuitable for interactive or time-sensitive applications such as AR or mobile capture.
  • (–) Requires expert setup: Users must collect dense, accurately posed input images to fit 3DGS properly, limiting accessibility for casual creators.
Single-pass methods (ours):
  • (+) Fast inference: The model runs in a single forward pass with no test-time optimization, enabling nearly real-time 3D stylization—ideal for interactive or mobile scenarios.
  • (+) User-friendly: The pipeline accepts a small set of casually captured images and produces a stylized 3D scene in an end-to-end manner, making it accessible to non-experts.
  • (–) Data-hungry: Training requires a large and diverse dataset to generalize across various scene types and artistic styles.
  • (–) Lower reconstruction quality: Feed-forward models typically prioritize speed and generalization over precise geometric fidelity.
  • (–) Limited scalability: Current implementations may face memory constraints (e.g., VRAM bottlenecks) and require architectural adjustments or more training data to handle large-scale environments.
@@ 6 Conclusion
Acknowledgements.
This work was supported in part by NSFC under Grant 62202389, in part by a grant from the Westlake University-Muyuan Joint Research Institute, and in part by the Westlake Education Foundation.
@@ Acknowledgements.
  • more visual results of our model and comparisons with baselines in Sec. B;
  • limitations of our model in Sec. C.
@@ Appendix B More Visual Results
Appendix C Limitations
Inherited from the base model [49], our method is currently limited to 2–8 input views and low-resolution scenarios ($256\times 256$). The former limitation could be alleviated by replacing the base model with VGGT [43], while the latter could be partially mitigated by fine-tuning the current model on higher-resolution data. However, ultra-high-resolution scenarios remain challenging due to VRAM constraints.